# Continuous and Discrete-Time Filters: A Unified Operational Perspective

### v1

@@ V-A1 Discrete low-pass filter
Then treating the initial condition as a discrete impulse $\delta[n]=x[0]=x_{0}$ is equivalent to considering a discrete system (Fig. LABEL:) with a transfer function between input and output given by:
@@ V-A2 Discrete high-pass filter
The stabilization of the IIR architecture is obtained by replacing the direct backward difference with a smoothed version, where the exponential weights are responsible for the averaging. The transfer function can be easily obtained by Z-transforming the recurrence of Eq- 65:
@@ V-C Sampled-Data Filter Implementations
Fig. 1 illustrates a FIR architecture implementing a discrete differentiator (ideal discrete derivation) using such an arrangement. The corresponding transient responses in Fig. 2 shows, as expected, a doublet for the impulse response and an ideal impulse for the step response. Mapping the discrete normalized frequency to its continuous-time equivalent in hertz yields the effective frequency response of a system that samples a CT signal, filters it in the discrete domain, and reconstructs it to continuous time. The equivalent CT response exhibits a maximum gain of 6dB6~dB at the Nyquist frequency $f=F_{s}/2$, while the zero at DC ($z^{-1}=1$) provides the expected attenuation of the low-frequencies.
Figure 1: Block diagram of a discrete differentiator with FIR architecture.
Figure 2: Transient (top) and frequency (bottom) response of the discrete differentiator.
@@
Fig. 3 depicts a FIR architecture implementing a two-sample moving sum. The transient responses in Fig. 4 demonstrates the effect of the two-units delay as the input impulse is broadened and the step’s rising edge is linearly interpolated. The equivalent CT frequency response, exhibits an all-pass filtering behavior with a gain of 6dB6~dB, due to the two-sample summation. The gain roll-off near the Nyquist frequency arises because successive samples span different half‑cycles of the input sinusoid. The resulting progressive phase shift leads to increasing cancellation in the summation, culminating in a complete null at the Nyquist point where $z^{.1}=-1$.
Figure 3: Block diagram of a 2-sample moving sum with FIR architecture.
Figure 4: Transient (top) and frequency (bottom) response of the 2-sample moving sum.
@@
Fig. 5 depicts an IIR architecture implementing an accumulator (ideal discrete integration). The transient responses in Fig. 6 illustrate the expected characteristic behavior by which an impulse is converted into a unit step and a step into a linear ramp. The equivalent CT frequency response reveals a pole responsible for infinite gain at DC. Whereas, near the Nyquist frequency the gain reaches a minimum of 6dB-6~dB reflecting the progressive cancellation due to phase misalignment. The pole reflects the unbounded growth of the step response. Ideally a discrete integrator accumulates without limit in practice, however, saturation constraints the operational range. Alternatively, free‑running accumulators can be employed within control loops, where they act as integrators whose unbounded drift is stabilized by negative feedback, effectively converting open‑loop integration into stable, closed‑loop integral sensing.
Figure 5: Block diagram of a discrete accumulator with IIR architecture.
Figure 6: Transient (top) and frequency (bottom) response of the discrete accumulator.
@@
Fig. 7 depicts an IIR architecture implementing a DT oscillator. The transient responses in Fig. 8 display a sustained oscillation at the Nyquist frequency. The CT equivalent frequency response reveals an unstable pole at $z=^{-1}=-1$, which produces infinite gain at that frequency while attenuating all others. Despite its intrinsic instability, this configuration is commonly employed as a clock‑frequency divider incorporating a sampled bi-stable element.
Figure 7: Block diagram of a discrete oscillator with IIR architecture.
Figure 8: Transient (top) and frequency (bottom) response of the discrete oscillator.
@@
The corresponding block diagram can be drawn from the discrete accumulator, Fig. 5, by incorporating the scaling factor associated with the pole $z_{p}$. The transient and frequency responses of Fig. 9 show the characteristic behavior of a leaky integration for frequencies well below the Nyquist point where the gain reaches its minimum $(1-z_{p})/(1+z_{p})$. This shows how the DT approximation only allows a limited attenuation of high-frequencies whose value is directly related to the pole responsible for the cut-off.
Figure 9: Transient (top) and frequency (bottom) of the discrete low-pass filter with $z_{p}=0.6$.
Unlike CT, where the sign of the real part controls growth/decay, in DT both positive and negative real poles inside the unit circle are stable, but a negative pole induces alternating sign oscillations. The limit case $z_{p}=-e^{i\pi}$ corresponds to the discrete oscillator depicted in Fig. 7; where the sign alternation arises because the impulse response consists of integer powers of the pole, as illustrated in Fig. 60. The pole’s phase induces an oscillation at the Nyquist frequency which can be interpreted as a train of discrete doublets. Inside the unit circle stability is insured by the damping action of the pole’s modulus, which imposes a decaying exponential envelope on the oscillation. Alternatively, the impulse response can be interpreted as a single doublet (typically realized with a differentiating positive zero) subjected to leaky accumulation by an IIR integrator. The overall impulse response, reproduced in Fig. 10, when convolved with an input sequence results in a leaky accumulation of past differences. Thus, a first-order system with a negative pole approximates a high‑pass filter by emulating a positive zero stabilized with a low‑pass pole; indeed, the resulting frequency response amplifies frequencies near the Nyquist point relative to lower ones, according to its transfer function:
@@
The corresponding block diagram can be obtained from the discrete oscillator of Fig. 7 by adding a proper scaling factor in the feedback path that moves the pole inside the unit circle. The derivative characteristic is also observable in the step response where the amplification of the rising edge induces an overshoot typical of underdamped systems which are designed near their stability boundary; in the continuous case this condition corresponds to a reduced phase margin.
Figure 10: Transient (top) and frequency (bottom) of a discrete first-order transfer function with negative zero $z_{p}=-0.6$.
Figure 11: Transient (top) and frequency (bottom) of the discrete high-pass filter with $z_{p}=0.6$.
@@
The corresponding block diagram can be obtained through the series connection of the differentiator, Fig, 1, and the accumulator, Fig, 5, with a proper scaling in the feedback to position the pole. This effectively adds a zero at DC to the low-pass filter of Eq. 76. The transient and frequency responses in Fig. 11 exhibit a well‑behaved high‑pass characteristic across the entire Nyquist band. In contrast, the corresponding low‑pass filter shows increasing deviation as the frequency approaches the Nyquist point, where aliasing and sampling effects become pronounced

### v2

@@ V-A1 Discrete low-pass filter
Then treating the initial condition as a discrete impulse $\delta[n]=x[0]=x_{0}$ is equivalent to considering a discrete system with a transfer function between input and output given by:
@@ V-A2 Discrete high-pass filter
The stabilization of the IIR architecture is obtained by replacing the direct backward difference with a smoothed version, where the exponential weights are responsible for the averaging. The transfer function can be easily obtained by Z-transforming the recurrence of Eq- 63:
@@ V-C Sampled-Data Filter Implementations
The FIR architecture Fig. 2-(a) with a unitary amplitude coefficient $z_{z}=1$ constitutes a discrete differentiator (ideal discrete derivation). The corresponding transient responses in Fig. 3 shows, as expected, a doublet for the impulse response and an ideal impulse for the step response. Mapping the discrete normalized frequency to its continuous-time equivalent in hertz yields the effective frequency response of a system that samples a CT signal, filters it in the discrete domain, and reconstructs it to continuous time. The equivalent CT response exhibits a maximum gain of 6dB6~dB at the Nyquist frequency $f=F_{s}/2$, while the zero at DC ($z^{-1}=1$) provides the expected attenuation of the low-frequencies.
Figure 2: Block diagram of basic first-order DT LTI systems constructed from unit delay, amplitude scaling, and summation elements. (a) FIR differentiator with zero at $z_{z}$. (b) FIR two-sample moving average with a negative zero at $z_{z}$. (c) IIR accumulator with pole at $z_{p}$. (d) IIR oscillator with a negative pole at $z_{p}$.
Figure 3: Transient (top) and frequency (bottom) response of the discrete differentiator. The impulse response is a discrete doublet which under convolution provides the difference between consecutive samples. The step response is a discrete impulse. The amplitude response shows a high-pass characteristic with a zero at DC $z=1$ and a maximum of 6dB6~dB at Nyquist z=1z=-1.
@@
Fig. 2-(b) depicts a FIR architecture implementing a two-sample moving sum. In the case of a zero at $z^{-1}=-1$, the transient responses shown in Fig. 4 demonstrates the effect of the two-units delay as the input impulse is broadened and the step’s rising edge is linearly interpolated. The equivalent CT frequency response, exhibits an all-pass filtering behavior with a gain of 6dB6~dB, due to the two-sample summation. The gain roll-off near the Nyquist frequency arises because successive samples span different half‑cycles of the input sinusoid. The resulting progressive phase shift leads to increasing cancellation in the summation, culminating in a complete null at the Nyquist point where $z^{-1}=-1$.
Figure 4: Transient (top) and frequency (bottom) response of the 2-sample moving sum. The impulse response is a 2-sample rectangular pulse which under convolution provides the sum between consecutive samples. The step response is a step with a linear rising edge. The amplitude response shows a pass-band with gain of 6dB6~dB from DC up to higher frequencies and a zero at Nyquist z=1z=-1.
@@
Fig. 2-(c) depicts an IIR architecture which with a unit pole reproduces an accumulator (ideal discrete integration). The transient responses in Fig. 5 illustrate the expected characteristic behavior by which an impulse is converted into a unit step and a step into a linear ramp. The equivalent CT frequency response reveals a pole responsible for infinite gain at DC. Whereas, near the Nyquist frequency the gain reaches a minimum of 6dB-6~dB reflecting the progressive cancellation due to phase misalignment. The pole reflects the unbounded growth of the step response. Ideally a discrete integrator accumulates without limit in practice, however, saturation constraints the operational range. Alternatively, free‑running accumulators can be employed within control loops, where they act as integrators whose unbounded drift is stabilized by negative feedback, effectively converting open‑loop integration into stable, closed‑loop integral sensing.
Figure 5: Transient (top) and frequency (bottom) response of the discrete accumulator. The impulse response is a step which is the integral of the impulse. Under convolution it produces accumulation with unitary weights of input samples. The step response is an unbounded linear ramp. The amplitude response shows a low-pass characteristic with a pole at DC $z=1$ and a minimum gain of 6dB-6~dB at Nyquist z=1z=-1.
@@
Fig. 2-(d) depicts an IIR architecture implementing a DT oscillator. The transient responses with a unit negative pole shown in Fig. 6 presents a sustained oscillation at the Nyquist frequency. The CT equivalent frequency response reveals an unstable pole at $z^{-1}=-1$, which produces infinite gain at that frequency while attenuating all others. Despite its intrinsic instability, this configuration is commonly employed as a clock‑frequency divider incorporating a sampled bi-stable element.
Figure 6: Transient (top) and frequency (bottom) response of the discrete oscillator. The impulse response is an oscillation at Nyquist which under convolution provides the sum of differences of pairs of consecutive samples. The step response is also an oscillation with smaller amplitude. The amplitude response shows a high-pass notch characteristic with a pole at Nyquist z=1z=-1 and a minimum gain of 6dB-6~dB away from Nyquist.
@@
The corresponding block diagram can be drawn from the discrete accumulator, Fig. 2-(c), by incorporating the scaling factor associated with the pole $z_{p}$. The transient and frequency responses of Fig. 7 show the characteristic behavior of a leaky integration for frequencies well below the Nyquist point where the gain reaches its minimum $(1-z_{p})/(1+z_{p})$. This shows how the DT approximation only allows a limited attenuation of high-frequencies whose value is directly related to the pole responsible for the cut-off.
Figure 7: Transient (top) and frequency (bottom) of the discrete low-pass filter with $z_{p}=0.6$. The impulse response is a low-passed impulse with the characteristic exponential decay. The step response is a step with exponential settling. The amplitude response shows a low-pass characteristic with a pole at $z_{p}$ and a minimum gain at Nyquist z=1z=-1.
Unlike CT, where the sign of the real part controls growth/decay, in DT both positive and negative real poles inside the unit circle are stable, but a negative pole induces alternating sign oscillations. The limit case $z_{p}=-e^{i\pi}$ corresponds to the discrete oscillator depicted inFig. 2-(d); where the sign alternation arises because the impulse response consists of integer powers of the pole, as illustrated in Fig. 58. The pole’s phase induces an oscillation at the Nyquist frequency which can be interpreted as a train of discrete doublets. Inside the unit circle stability is insured by the damping action of the pole’s modulus, which imposes a decaying exponential envelope on the oscillation. Alternatively, the impulse response can be interpreted as a single doublet (typically realized with a differentiating positive zero) subjected to leaky accumulation by an IIR integrator. The overall impulse response, reproduced in Fig. 8, when convolved with an input sequence results in a leaky accumulation of past differences. Thus, a first-order system with a negative pole approximates a high‑pass filter by emulating a positive zero stabilized with a low‑pass pole; indeed, the resulting frequency response amplifies frequencies near the Nyquist point relative to lower ones, according to its transfer function:
@@
The corresponding block diagram can be obtained from the discrete oscillator of Fig. 2-(d) by adding a proper scaling factor in the feedback path that moves the pole inside the unit circle. The derivative characteristic is also observable in the step response where the amplification of the rising edge induces an overshoot typical of underdamped systems which are designed near their stability boundary; in the continuous case this condition corresponds to a reduced phase margin.
Figure 8: Transient (top) and frequency (bottom) of a discrete first-order transfer function with negative zero $z_{p}=-0.6$. The impulse response is an oscillation at Nyquist with an exponential decay envelope. The step response is an underdamped step with exponential settling. The amplitude response shows a gradual high-pass notch characteristic with a negative pole at $z_{p}$ a maximum gain at Nyquist z=1z=-1.
Figure 9: Transient (top) and frequency (bottom) of the discrete high-pass filter with $z_{p}=0.6$. The impulse response is a high-passed impulse with the characteristic exponential recovery. The step response is an impulse with exponential decay. The amplitude response shows a high-pass characteristic with a pole at $z_{p}$ and a maximum gain at Nyquist z=1z=-1.
@@
The corresponding block diagram can be obtained through the series connection of the differentiator, Fig. 2-(a), and the accumulator, Fig. 2-(c), with a proper scaling in the feedback to position the pole. This effectively adds a zero at DC to the low-pass filter of Eq. 74. The transient and frequency responses in Fig. 9 exhibit a well‑behaved high‑pass characteristic across the entire Nyquist band. In contrast, the corresponding low‑pass filter shows increasing deviation as the frequency approaches the Nyquist point, where aliasing and sampling effects become pronounced